### v1

# Against ‘softmaxing’ culture

Abstract
AI is flattening culture. Evaluations of “culture” are showing the myriad ways in which large AI models are homogenizing language and culture, averaging out rich linguistic differences into generic expressions. I call this phenomenon “softmaxing culture,” and it is one of the fundamental challenges facing AI evaluations today. Efforts to improve and strengthen evaluations of culture are central to the project of cultural alignment in large AI systems. This position paper argues that machine learning (ML) and human-computer interaction (HCI) approaches to evaluation are limited. I propose two key shifts. First, instead of asking “what is culture?” at the start of system evaluations, I propose beginning with the question: “when is culture?” Second, while I acknowledge the philosophical claim that cultural universals exist, the challenge is not simply to describe them, but to situate them in relation to their particulars. Taken together, these conceptual shifts invite evaluation approaches that move beyond technical requirements, toward perspectives more responsive to the complexities of culture.
keywords:
Culture, sociolinguistics, low-resource languages, representation, stereotypes, discourse, dialogue and pragmatics, ChatGPT, NLP, Africa
1 Introduction
Large AI models, particularly large language models (LLMs), mediate how culture is represented, interpreted, and reproduced (Liu et al., 2025; Adilazuarda et al., 2024; Zhou et al., 2025; Bajohr, 2025; Farrell et al., 2025). These systems promise universal fluency and translation across linguistic and cultural boundaries, but in doing so, they often flatten the complexity of culture into generic forms. I refer to this phenomenon as “softmaxing culture.” [^1] In ML, the softmax function transforms a set of raw output scores into a probability distribution, amplifying the most likely outputs while reducing the less probable. This probabilistic logic, when applied to culture, has a clear effect: it privileges dominant, statistically frequent expressions while suppressing non-frequent expressions. What is statistically probable is often confused with what is socially or culturally appropriate. It is also important to note that the phenomenon I have described has nothing to do with ‘softmaxxing,’ a beauty trend associated with enhancing one’s appearance. A rapidly expanding body of work is showing how LLMs are driving diverse users of English as a second language to Western linguistic norms and patterns (Bommasani et al., 2022; Agarwal et al., 2025); how LLMs propagate racial, geographic, and gender-based stereotypes (Cheng et al., 2023; Decoupes et al., 2025); and judgments by LLMs are prejudiced in scoring comparisons around diversity (Ye et al., 2024). While NLP is grappling with the challenges of these large AI models, including the biases and stereotypes that these models propagate (Bender et al., 2021), the evaluation of these systems is limited (Wallach et al., 2025). This limitation stems from the fact that model outputs are statistical predictions of patterns learned from large corpora of training data. As such, system evaluations tend to measure narrowly defined aspects of language and culture, leaving out the more complex, relational content that constitutes culture and language.
Despite differences in method and practice between ML and HCI research, [^2] It is true, ML research is re-inventing HCI’s core problematic – user research. However, when it comes to the “softmaxing culture” problem, neither community seems to offer satisfying answers. HCI tends to suggest that ML does not adequately understand human-centered research; meanwhile, ML is quick to point out HCI’s technical gaps. researchers across these fields are drawing on established social science criteria for evaluation (Hofstede, 2001, 2011), as well as constructing benchmarks informed by varied interpretations of language and culture (Shi et al., 2024; Adelani et al., 2025; Singh et al., 2025; Masoud et al., 2024; Ochieng et al., 2024). Indeed, across recent survey publications, the majority of papers examined in ML and NLP are drawing on methods such as Hofstede’s dimensions of culture and the World Values Survey. These methods have been instructive in generating useful and actionable insights into the workings of large AI models. While the ML and HCI are adopting socio-technical approaches towards culturally aware models, there are limitations. These approaches are plagued by fundamental gaps in social and cultural understanding of the values and assumptions embedded in the data and benchmarks leveraged in the evaluation of large AI models. This position paper presents two conceptual shifts: first, the question “what is culture?” is the wrong question to start with in scoping and operationalizing culture in LLMs; second, following Wiredu (1995)’s claim that there are cultural universals, the more pragmatic question to ask is “when is culture?” The goal lies in understanding relations between universals and particulars, and when they become relationally valid and tenable. While there are growing demands emerging for context-driven and more situated evaluations beyond the artifact into real-world contexts, for the sake of this paper, I will focus on the conceptual aspects of culture and its formulation in evaluation. By re-orienting our understanding of culture, we will have a better understanding of assessing cultural values and constructs in large AI models at different stages of development and use.
2 Evaluating ‘culture’: How ML and HCI have a lot in common than they acknowledge
Both ML and HCI recognize that context matters, albeit differently. On one hand, ML research (since the deep learning revolution) has been preoccupied with advancing the frontiers of AI capabilities by building on the “unreasonable effectiveness” of big data (Halevy et al., 2009). These advancements have been complemented by the rapid development and adoption of benchmarks to measure and compare modes across tasks and capabilities. Breakthroughs in ML, particularly modern approaches to embeddings, [^3] Notably neural network-based approaches including word2vec, seq2seq, and transformers trained on enormous amounts of data. showed meaningful connections between words and sentences, providing the capacity for modeling and measuring cultural contexts through large AI models. On the other hand, HCI has been engaged with understanding users in the contexts in which they interact with computing technologies. Sensitized in some part by the generative model of culture, HCI examines how relationships between users and “technological objects and knowledge practices of everyday life become meaningful contingently and dynamically as social activity unfolds” (Irani et al., 2010).
But what is hardly surprising is that a growing body of recent work in ML and HCI shows limitations in the framing and operationalization of culture in system evaluations (Liu et al., 2025; Adilazuarda et al., 2024; Zhou et al., 2025). Whereas there is increasing recognition across the ML and HCI divide that culture is dynamic and relational, most evaluations still appear to capture narrow definitions and static snapshots of “culture.”
3 What is culture? – and why it is the wrong question to begin with
While there are several definitions of culture, they are out of scope of this paper. I want to focus on Hofstede’s definition and taxonomy of culture that have achieved wide acceptance and acclaim in the social sciences since the 1970s and more recently in ML evaluation research. According to Hofstede (2011), “culture is the collective programming of the mind that distinguishes the members of one group or category of people from others.” Culture is software for the society of minds. Hofstede (2001, 2011) across several decades of work has developed a leading framework for evaluating culture: Hofstede’s Dimensions of Culture (Hofstede, 2011). Hofstede’s cultural dimensions describe how societies – countries and institutions – vary in their values and practices, and these dimensions are categorized in six ways: power distance (acceptance of hierarchy); uncertainty avoidance (comfort with ambiguity); individualism versus collectivism (self vs. group); masculinity versus femininity (competition vs. care); long-term versus short-term orientation (future vs.tradition), indulgence versus restraint (freedoms vs. control).
NLP and sociolinguistics scholars Zhou et al. (2025) have provided one of the most compelling cases against mainstream evaluations of culture in AI, suggesting that macro-level definitions of culture significantly constrain how culture is understood and evaluated. These definitions are unsatisfactory. Instead the scholars argue for “cultural competence” for NLP, with renewed attention to localization. Many grassroots efforts such as Masakhane and Deep Learning Indaba have made considerable progress and continue to show great promise on this path of localizing and contextualizing language models (Nekoto et al., 2020). Indeed benchmarks and fine-tuning, often constructed from curations of data workers, Wikipedia articles, and corpora of community-specific social media posts, have advanced our understanding of cultural alignment in AI. However, culture is not about the identification of particular symbols, ideologies, cognitive patterns, etc. Culture is dynamic and complex. As such, attempting to nail down culture to a narrow definition – captured in datasets – only risks stereotyping and essentializing certain people (and their cultures) (Zhou et al., 2025; Bender et al., 2021). The question “what is culture [in AI]?” is therefore the wrong question to start with. The more productive question to ask, therefore, is “when is culture?” This question challenges researchers to fully examine what it means for culture to be relationally valid and tenable. Implicit in this question, too, is the partiality of our understanding of culture – and how this must be accounted for in evaluations.
4 Are there cultural universals?
The short answer is yes. For Wiredu (1995, 1996), a Ghanaian philosopher, if there were no cultural universals, there would be no intercultural communication. Cultural universals – notably through language (an artifact of culture) – enable the necessary translation work. Although his argument is simple, it is relevant to evaluations of culture in AI in at least two ways. First, humans generally share a ‘cognitive toolkit’ that endows them with capabilities such as reflective perception (possession of a concept of the external world), abstraction (bringing particulars under general concepts), deduction (applying rules to observations), and induction (learning from examples). [^4] While the deduction can be made that large AI models are also endowed a related ‘cognitive toolkit,’ those arguments are beyond the scope of this paper. Second, and importantly, humans possess the principle of sympathetic impartiality; a certain due concern for others. Now our role here, as set up in the first point, is to identify the universals – and of course the particulars. To be sure, this point is not about definitions, but rather demands that we think carefully about the human project; what makes us human, irrespective of our geographic, historical, moralistic, and even linguistic differences, and how we can account for these differences in our research.
5 Beyond ‘softmaxing’ culture: A more pragmatic research vision
The use of the metaphor softmaxing culture is closely related to the tendencies of LLMs to homogenize text and media. These tendencies are the feature of transformer-based large AI systems, and they have been exacerbated cultural incongruencies of AI (Prabhakaran et al., 2022): the odds between the cultural assumptions baked into systems (both explicitly and from data sources) and the expectations of the target cultural ecosystems these technologies are used. Despite these structural issues, I argue that researchers in ML and HCI working around issues and questions of culture should focus on understanding the relational contexts through which cultural values and practices are made intelligible, especially for low-resource languages and expressions that are underrepresented in pre-training and fine-tuning datasets. Going local is an early sign of success. However, confronting the softmaxing problem will entail deep and meaningful integration of theoretical and empirical strategies for studying cultures beyond the ML/HCI divide.